Ricardo is 4 years less than twice as old as his sister. The sum of their ages is 20. How old is Ricardo?
Answer by checkley71(8403)

R+S=20
R=2S-4 THUS
(2S-4)+S=20
2S-4+S=20
3S=20+4
3S=24
S=24/3
S=8 THE R=2*8-4 OR R=16-4 OR R=12
PROOF
12+8=20
20=20